Show the effect of pandemic an piyush optimal consumption Bundy assuming that substitution effects outweighs the income effects for book

1
Expert's answer
2022-02-16T10:40:19-0500

Piyush budget constraint before pandemic was:

"(PF1\u00d7QF)+(PB\u00d7QB)="

"Y(PF1\u00d7QF1)+(PB\u00d7QB)=Y"

Where:

PF=price of food.

QB= quantity of books.

QF= quantity of food.

PB= Price of books.

Substitution effect will increase the quantity demanded for books from QB to QB1 thus the new budget constraint will be as follows:

"(PF1\u00d7QF1)+(PB\u00d7QB1)="

"Y(PF1\u00d7QF1)+(QB1\u00d7QB)=Y"